OpenStudy (anonymous):

Can someone please help, Simplify:

OpenStudy (anonymous):

\[\sqrt[3]{54(x ^{4}y ^{2})}\]

OpenStudy (anonymous):

hmm it says the answer is: \[3x\sqrt[3]{2xy ^{2}}\]

OpenStudy (anonymous):

Im just confused on the steps :S

OpenStudy (amistre64):

with a cube root, you want to find things under it that have a multiplicity of 3 1.1.1 = 1 2.2.2 = 8 3.3.3 = 27 4.4.4 = 64 5.5.5 = 125 etcc

OpenStudy (amistre64):

these can be factored, or pulled out

OpenStudy (amistre64):

the variable can be "seen" since their exponents already tell us their multiplicity

OpenStudy (anonymous):

Because of cube root, you try to rewrite expression under the radical in cubes if possible. So, 54x^4y^2=2*27*x^3*x*y^2=2*3^3*x^3*y^2 Then you can take cube root of 3^3 and x^3 3x cube root(2x*y^2)

OpenStudy (anonymous):

Oh ok, thanks a lot for your help everyone.

OpenStudy (anonymous):

You are welcome. You always can check with this website for answers or at least ideas http://www.wolframalpha.com/input/?i=%2854x%5E4+y%5E2+%29%5E%281%2F+3%29++